OpenStudy (anonymous):

What are the exact solutions of x2 + 5x + 2 = 0? A. x= -5+/- square root 33/2 B. x=5 +/- square root 33/2 C. x= -5+/- square root 17/2 D. x= 5+/- square root 17/2

OpenStudy (anonymous):

C

OpenStudy (anonymous):

The answer is B.

OpenStudy (anonymous):

\[x^2 + 5x + 2 = 0\]\[x^2 + 5x =-2\]\[x^2 + 5x+25/4 =-2+25/4\]\[(x+5/2)^2=17/4\]\[x+5/2=\pm \sqrt{17}/2\]\[x=-5/2 \pm \sqrt{17}/2=\frac{-5 \pm \sqrt {17}}{2}\]